# On Detecting Cherry-picking in News Coverage Using Large Language Models

### v1

@@ Abstract
Cherry-picking refers to the deliberate selection of evidence or facts that favor a particular viewpoint while ignoring or distorting evidence that supports an opposing perspective. Manually identifying instances of cherry-picked statements in news stories can be challenging, particularly when the opposing viewpoint’s story is absent. This study introduces Cherry, an innovative approach for automatically detecting cherry-picked statements in news articles by finding missing important statements in the target news story. Cherry utilizes the analysis of news coverage from multiple sources to identify instances of cherry-picking. Our approach relies on language models that consider contextual information from other news sources to classify statements based on their importance to the event covered in the target news story. Furthermore, this research introduces a novel dataset specifically designed for cherry-picking detection, which was used to train and evaluate the performance of the models. Our best performing model achieves an F-1 score of about 89% in detecting important statements when tested on unseen set of news stories. Moreover, results show the importance incorporating external knowledge from alternative unbiased narratives when assessing a statement importance.
@@ 1 Introduction
The erosion of citizens’ trust in the media has reached a critical stage, posing significant harm to the foundations of democracy Dan et al. (2021); Druckman and Parkin (2005); Chiang and Knight (2011). Research shows evidence indicating a decline in media trust among citizens  Strömbäck et al. (2020). Media losing trust can be attributed to multiple factors, among which, two important factors are the perception of political bias in media reporting Lee (2010) and consumers’ exposure to misinformation and disinformation Lee et al. (2023); Hameleers et al. (2022).
@@
Cherry-picking, a widely used but often imperceptible technique in news reporting, is a device that aids bias and disinformation Vincent (2019); Paul and Elder (2019). It is generally defined as a logical fallacy that involves deliberately selecting (and/or exaggerating) facts that bolster a certain argument and omitting (and/or under-emphasizing) facts that support the opposing viewpoint Best (2004). Common aliases for cherry-picking known in literature are one-sideness Paul and Elder (2008), argument by half-truth Sproule (2001), and case-making Lee (1945).
@@
Cherry-picking occurs in a variety of domains, such as media narratives, science and politics Best (2004). Cherry-picking can be observed across scales as large as in selecting which events to cover, and as subtle as in choosing specific words for a particular story. Furthermore, cherry-picking can manifest in various types of data. For example, for facts staffed with numbers and stats, cherry-picking can occur when specific segments of a trend-line Asudeh et al. (2020) or particular statistics Best (2004) are chosen to bolster a particular claim Wu et al. (2017). For facts that are visual (e.g., videos), cherry-picking occurs when only certain frames are selectively presented.
@@
The focus of this study is cherry-picking in statements within news articles, particularly when only certain facts are included while others are taken out. A multitude of media outlets, including prominent and highly regarded ones, exhibit biases in various topics within domains such as politics and science Baron (2006). These media outlets often exercise caution in curating the facts included in their stories covering controversial topics Paul and Elder (2019, 2008). This selection of facts is capable of swaying the public opinion on crucial matters, such as climate change, vaccination, or elections. It can mislead readers into believing that the partial truth presented in these stories represents the entire truth Fleming (1995).
We introduce a novel method for automatically detecting cherry-picking in news reports by leveraging context derived from the covered event. Our method employs language models, including BERT and Longformer, trained to assess a statement’s importance based on context. The models are trained on the novel Cherry dataset we curated for this purpose. The proposed method shows promising results in predicting a statement’s importance given external contextual information. Our top-performing model achieves about 89% in terms of F-1 score and about 90% in terms of accuracy. To the best of our knowledge, this is the first study that addresses the detection of cherry-picking in text using computational approach. The key contributions of this work are as follows: (a) The formulation and modeling of the problem from a computational perspective. (b) The novel end-to-end cherry-picking detection approach that incorporates contextual information. (c) Additionally, we introduce a new cherry-picking detection dataset. All artifacts related to this work can be found at our anonymous Github repository https://github.com/emnlp-cherry/Cherry.
@@ 2 Related Work
The research on media bias can be classified into three levels according to text granularity: word-level, sentence-level, and article-level Chen et al. (2020). Nevertheless, distinguishing between these three levels is not always straightforward, as they often intersect and influence each other to a certain degree. For instance, word-level bias is dependent on sentence or article context, whereas sentence-level bias relies on the broader article context.
@@
•  Word-level bias. The choice of words in news articles can influence the audience’s opinions towards a particular event Hamborg (2020). Word-level bias encompasses two primary categories: framing bias and epistemological bias Recasens et al. (2013). The connotations associated with the same words can vary based on the specific context in which they are employed Spinde et al. (2021b).
@@
•  Sentence-level bias. Numerous studies focus on detecting bias at sentence level, considering it as the fundamental level that can be combined with article-level bias Spinde et al. (2021a). The study in Hube and Fetahu (2018) introduced an automated method for identifying biased statements from Wikipedia using lexicon of bias words. In Lei et al. (2022), the authors studied the task of identifying biased sentences in news articles that serve to explain the overall bias present throughout the whole article. Focusing on media’s sentiment toward target entities, Fan et al. (2019) examined sentence-level bias along with article context.
@@
•  Article-level bias. For identifying political ideology in news articles, Baly et al. (2020) constructed a dataset of news articles annotated with left, center, and right leanings, and built and evaluated machine learning models using the dataset to predict the articles’ ideological leaning.
AllSides.com, AdFontesMedia.com, and MBFC are widely utilized sources in media bias analysis, providing researchers with tools and ratings to evaluate the political bias and credibility of news sources. They employ diverse methods, such as community feedback, analysts with varying viewpoints, and a combination of objective measures and subjective analysis, to evaluate bias in news outlets. Literature addressing bias and misinformation detection heavily relies on these three sources for bias categorization in tasks such as source annotation on social media platforms Weld et al. (2021), comparing the diffusion of news from reliable and non-reliable sources Cinelli et al. (2020), benchmarking automatic media bias monitors Ribeiro et al. (2018), and creating corpora for automatic media bias detection Chen et al. (2020).
While previous media bias research primarily focuses on predisposition towards a particular viewpoint, such as ideological leanings and sentiment bias, there has been limited research on the intentional practice of cherry-picking in news articles, which is a tactic employed to selectively present information that supports a certain bias. Although existing studies have explored the impact of cherry-picking in various contexts, such as computational fact-checking Wu et al. (2017); Lin et al. (2021) and detecting cherry-picking in trendlines Asudeh et al. (2020), they have not specifically addressed cherry-picking in news articles. While Ehrhardt et al. (2021) introduced a method to predict news articles’ ideological leaning by examining the omission of entities, our approach distinguishes itself by focusing on detecting cherry-picked statements in news articles, instead of analyzing co-occurrences of entities across different articles.
Diverging from previous research, our study stands out as the first comprehensive examination of cherry-picking in news articles, specifically analyzing the importance of statements extracted from news articles and analyze whether these statements are underemphasized by the articles, thus constituting instances of cherry-picking. In light of this, we have undertaken the construction of a new cherry-picking dataset, providing a valuable resource for evaluating the prevalence and impact of cherry-picked statements in news articles. Moreover, we developed a cherry-picking detection pipeline to identify cherry-picked statements. Our contribution could shed light on the manipulation of information in news articles. It not only enhances our understanding of mechanisms behind cherry-picking but also contributes to the promotion of more objective and informed journalism.

### v2

# On Context-aware Detection of Cherry-picking in News Reporting

@@ Abstract
Cherry-picking refers to the deliberate selection of evidence or facts that favor a particular viewpoint while ignoring or distorting evidence that supports an opposing perspective. Manually identifying cherry-picked statements in news stories can be challenging. In this study, we introduce a novel approach to detecting cherry-picked statements by identifying missing important statements in a target news story using language models and contextual information from other news sources. Furthermore, this research introduces a novel dataset specifically designed for training and evaluating cherry-picking detection models. Our best performing model achieves an F-1 score of about 89% in detecting important statements. Moreover, results show the effectiveness of incorporating external knowledge from alternative narratives when assessing statement importance.
@@ 1 Introduction
The erosion of citizens’ trust in the media has reached a critical stage, posing significant harm to the foundations of democracy Dan et al. (2021); Druckman and Parkin (2005); Chiang and Knight (2011). Research shows evidence of a decline in media trust among citizens  Strömbäck et al. (2020). Media losing trust can be attributed to the perception of political bias in media reporting Lee (2010) and consumers’ exposure to disinformation Lee et al. (2023); Hameleers et al. (2022).
@@
Cherry-picking, a widely used but often imperceptible practice in news reporting, is a device that aids bias and disinformation Vincent (2019); Paul and Elder (2019). It is defined as a logical fallacy that involves deliberately selecting (or exaggerating) facts that bolster a certain argument and omitting (or under-emphasizing) facts that support the opposing viewpoint Best (2004). Its common aliases in literature include one-sideness Paul and Elder (2008), argument by half-truth Sproule (2001), and case-making Lee (1945).
@@
In news reporting, cherry-picking can be observed across scales as large as in selecting which events to cover, and as subtle as in choosing specific words for a particular story. Furthermore, cherry-picking can manifest in various types of data such as charts  Asudeh et al. (2020) or when particular statistics Best (2004) are chosen to bolster a certain claim Wu et al. (2017).
@@
This study’s focus is cherry-picking statements within news articles, particularly when only certain statements are included while others are omitted. A multitude of media outlets, including prominent and highly regarded ones, exhibit biases in various topics within domains such as politics and science Baron (2006). These media outlets often exercise caution in curating the statements included in their stories covering controversial topics Paul and Elder (2019, 2008). This selection of statements is capable of influencing public opinion on pivotal subjects, such as climate change, vaccination, or elections, by misleading readers to perceive partial truth as holistic truth Fleming (1995).
We introduce a novel method for identifying cherry-picking in news stories by leveraging context derived from other narratives. Our method employs language models, including fine-tuned embedding models and few/zero shot prompting of generative models, to assess statements’ importance based on a given context to identify cherry-picked ones. The fine-tuned models are trained on the novel Cherry dataset we curated. The proposed method shows promising results in inferring a statement’s importance given external contextual information. Our top-performing model achieves an F-1 score of 89% and an accuracy of 90%.
To the best of our knowledge, this is the first study that addresses the detection of cherry-picking in text using computational approach. The key contributions of this work are as follows: (a) The formulation and modeling of the problem from a computational perspective. (b) The novel end-to-end cherry-picking detection approach that infuses contextual information with statements through different language modeling techniques. (c) A new cherry-picking annotated dataset which can prove useful in not only detection but also other lines of research related to cherry-picking. (d) Thorough experimentation to determine the most effective context to consider when assessing cherry-picking. All artifacts of this work are released under the GNU General Public License v3.0 at our anonymous Github repository https://github.com/cherry-pic/Cherry.
@@ 2 Related Work
Research on media bias can be classified into three levels based on text granularity: word-level, sentence-level, and article-level Chen et al. (2020). Distinguishing between these levels is not always straightforward, as they often intersect and influence each other. For instance, word-level bias is dependent on sentence or article context, and similarly sentence-level bias is contingent on the broader article context.
@@
•  Word-level bias. The choice of words in news articles can influence the audience’s opinions towards a particular event Hamborg (2020). Word-level bias encompasses two primary categories: framing bias and epistemological bias Recasens et al. (2013). The connotations associated with the same word can vary by the specific context in which they are employed Spinde et al. (2021b).
@@
•  Sentence-level bias. Numerous studies focus on detecting bias at sentence level, considering it as the fundamental level that can be combined with article-level bias Spinde et al. (2021a). Using lexicon of bias words, Hube and Fetahu (2018) introduced a method for identifying biased statements from Wikipedia. In Lei et al. (2022), biased sentences were identified in news articles, serving to explain the overall bias present throughout articles. Focusing on media’s sentiment toward target entities, Fan et al. (2019) examined sentence-level bias along with article context.
@@
•  Article-level bias. Baly et al. (2020) constructed a dataset of news articles annotated with left, center, and right leanings, then used machine learning models to predict articles’ political ideology.
Prior research on media bias primarily focused on predisposition towards particular viewpoints, leaving a notable gap in exploring media bias related to cherry-picking in news articles. Some studies have examined the impact of cherry-picking in computational fact-checking Wu et al. (2017); Lin et al. (2021) and the detection of cherry-picking in trendlines Asudeh et al. (2020). Nevertheless, there has been no specific investigation into end-to-end cherry-picking detection in news articles.